AMS Special Session on Geometry of Curves and Surfaces / Twelfth Annual Southeast Geometry Conference,

March 16-18, 2001.

The special session represents this year’s incarnation of the annual SouthEast Geometry Conference. It will be held at the University of South Carolina, Columbia SC on Friday, Saturday, and Sunday, March 16-18, 2001.

SEGC funds are available to those without grant support who wish to attend the conference/special session. Preference will be given to graduate students and other young mathematicians. It is not necessary to contribute a talk to take advantage of these funds. For more information contact Joe

Fu at fu@math.uga.edu.

The full program for the AMS meeting can be found at

http://www.ams.org/amsmtgs/2048 special.html.

Information about registration, accommodations, and the like can be found at

http://www.ams.org/amsmtgs/2048 other.html.

More information about the SouthEast Geometry Conference is at

http://www.math.sc.edu/∼howard/SEGC/.

Friday, March 16, Afternoon Session:

2:30–2:50 Thomas A. Ivey  Integrable Geometric Evolution Equations for Curves.
3:00–3:20 Alina Stancu  Evolving piecewise linear convex curves in the plane.
3:30–3:50 Eddie Fuller  Constructing Knot Isotopies with Geometric Constraints

Using Holonomic Space Curves.

4:00–4:20 Joseph H. Fu  A notion of criticality for ropelength.
4:30–4:50 Rob Kusner  Ropelength, thickness and conformal modulus of links.
5:00–5:45 Jason H. Cantarella Many knot energies are minimized by circles.

Saturday, March 17, Morning Session:

8:30–8:50 Mohammad Ghomi Skew Loops and Quadric Surfaces.
9:00–9:20 Henry C. Wente  Explicit Solutions to the H-Surface Equation on Tori.
9:30–9:50 Stephanie Alexander The Convex Hull Property and Topology of Hypersurfaces

with Nonnegative Curvature.

10:00–10:45 John M. Sullivan  The Second Hull of a Knotted Curve.

Saturday, March 17, Afternoon Session:

3:00–3:20 Andrejs E. Treibergs Nonpositively curved surfaces in Euclidean Space.
3:30–3:50 Joel L. Weiner  Isometric immersions of the Euclidean plane into Euclidean

4-space.

4:00–4:20 Scott D. Pauls  Smooth minimal surfaces in the Heisenberg group.
4:30–4:50 Jie Qing  On the finite time singularities of the heat flow for harmonic

maps from surfaces.

5:00–5:20 John McCuan  Continua of singular minimal graphs.
5:30–6:15 Frank Morgan  Area-minimizing Surfaces in Cones and Double Bubbles.

Sunday, March 18, Morning Session:

9:30–9:50 Ralph Howard  Unfoldings of Space Curves.
10:00–10:20 Shihshu Walter Wei Minimal Varieties and p-harmonic maps.
10:30–10:50 Elisabeth M. Werner An Analysis of Completely-Positive Trace-Preserving Maps on M_{2}.
11:00–11:20 Richard L. Bishop  Spines and Topology of Thin Riemannian 3-Manifolds.
11:30–11:50 David A. Singer  Diffeomorphisms of the Circle and Hyperbolic Curvature.
12:00–12:20 Svetlana A. Krat  On pairs of metrics invariant under a co-compact action of a group .

For more information contact one of the orignizers: Mohammad Ghomi ghomi@math.sc.edu or Ralph

Howard howard@math.sc.edu